### v3

# Bridging Quantum Mechanics to Organic Liquid Properties via a Universal Force Field

Abstract
Molecular dynamics (MD) simulations are essential tools for unraveling atomistic insights into the structure and dynamics of condensed-phase systems. However, the universal and accurate prediction of macroscopic properties based on ab initio calculations remains a significant challenge, often hindered by the trade-off between computational cost and simulation accuracy. Here, we present ByteFF-Pol, a graph neural network (GNN)-parameterized polarizable force field, trained exclusively on high-level quantum mechanics (QM) data. Leveraging physically-motivated force field forms and training strategies, ByteFF-Pol exhibits exceptional performance in predicting thermodynamic and transport properties for a wide range of small-molecule liquids and electrolytes, outperforming state-of-the-art (SOTA) classical and machine learning force fields. The zero-shot prediction capability of ByteFF-Pol bridges the gap between microscopic QM calculations and macroscopic liquid properties, enabling the exploration of previously intractable chemical spaces. This advancement holds transformative potential for applications such as electrolyte design and custom-tailored solvent, representing a pivotal step toward data-driven materials discovery.
[…]
Tianze Zheng at , Wen Yan at
Molecular dynamics (MD) simulations have emerged as one of the most important cornerstones of modern materials and biological research, offering an indispensable tool for probing condensed-phase systems at the atomic level [29, 20, 53]. By revealing microscopic details of complex phenomena such as structural rearrangements and molecular diffusion—often at resolutions inaccessible to experiments—MD simulations provide critical insights for applications ranging from drug discovery to material engineering [41, 47, 1]. Among these applications, MD simulations play a vital role in the screening and design of organic molecules, such as electrolytes for novel batteries [16, 17] and deep eutectic solvents for green chemistry [50]. They enable the rapid exploration of vast chemical and formula spaces for organic materials, which is impossible to enumerate through experiments.
The accuracy of these simulations, however, is critically dependent on the force field—a mathematical model describing interatomic interactions. This dependency necessitates the development of universal force fields combining accuracy and efficiency. Traditional force fields, such as Amber [22], CHARMM [7], and OPLS [26, 25], model interatomic interactions using a set of simple and predefined functional forms and tabulated parameters. Conventional wisdom indicates that the accuracy of traditional force fields for macroscopic properties essentially relies on error cancellation at microscopic level, a consequence of their limited expressive power. To achieve this cancellation, the force field parameters derived from low-level quantum mechanics (QM) calculations need to be further optimized by experimental data, including spectroscopic measurements [19], and thermodynamic properties such as densities and evaporation enthalpies [26, 25].
In contrast, machine learning (ML)-based force fields utilize the powerful fitting capabilities of neural networks to directly learn interatomic interactions from QM calculations, thereby achieving high accuracy in predicting intramolecular energy landscapes, spectroscopic properties, and related phenomena [31, 32, 27, 39]. However, comparing to traditional force fields, ML force fields without physical constraints require tremendous QM data for training. The development of a universal many-body ML force field is exceptionally challenging due to the vast chemical space of organic molecules and the daunting computational cost of generating sufficient ab initio data. Consequently, prominent ML force fields like MACE-OFF can suffer from accuracy and transferability issues, resulting in inferior performance in bulk property predictions when compared to well-established traditional ones [32]. While other efforts, such as the BAMBOO force field  [17], have demonstrated state-of-the-art (SOTA) performance in predicting the bulk properties of electrolytes, they still depend on experimental density data for empirical fine-tuning [17]. Ultimately, the predictive capability of both traditional and ML force fields remains constrained by the availability and integration of experimental information.
To address this challenge, efforts have been made toward establishing ab initio force fields by refining traditional functional forms to better incorporate the underlying physics of molecular interactions [52, 23]. These force fields are typically based on separating the intermolecular interactions into distinct components [43, liu2019amoeba+] or employing many-body expansion [2]. Despite the solid physical foundation, such force fields often struggle with computational efficiency and transferability [52]. Polarizable force fields, including AMOEBA [40, 9, liu2019amoeba+], APPLE&P [4] and CL&Pol [16], have emerged as a successful framework that retains computational efficiency while capturing the response of electron density to different electrical environments—a critical feature for systems like electrolytes [30]. Nevertheless, parameterizing these polarizable force fields remains a complex, labor-intensive process that often relies on manual adjustments. As a result, their transferability across different chemical environments is uncertain, failing to consistently outperform traditional force fields when applied to systems with varying electronic structures or bonding characteristics.
In summary, currently there is still an urgent need for an universal organic force field that possesses both the accuracy of high-level ab initio calculations, as well as running speed, generalizability, and data efficiency required by large-scale simulations of complex materials. In response to this challenge, we introduce ByteFF-Pol, a graph neural network [10] (GNN)-parameterized polarizable force field that represents a significant advancement in force field development. Unlike traditional and most ML-based force fields, ByteFF-Pol is trained exclusively on high-level QM labels, eliminating the need for experimental calibration. Through a novel force field architecture and training scheme, ByteFF-Pol achieves exceptional accuracy in predicting the thermodynamic and transport properties of small-molecule liquids and electrolytes, outperforming SOTA traditonal and ML force fields. Its zero-shot prediction capabilities enable the exploration of previously intractable chemical spaces, opening up new avenues for data-driven discovery of materials with optimized properties for applications in electrolyte design, complex solvents, and beyond.
1 Results
1.1 Overview of ByteFF-Pol
The framework of ByteFF-Pol is illustrated in Figure 1. Unlike traditional methods that rely on look-up tables, ByteFF-Pol employs a GNN model to predict force field parameters directly from molecular graphs. During the training stage, these parameters are integrated with atomic coordinates into energy functions to compute the decomposed energy terms, which are then fitted to reference labels from the absolutely localized molecular orbital energy decomposition analysis (ALMO-EDA) calculations [28]. After training, the parameters of the GNN model are fixed. The force field parameters predicted by the GNN model can be directly used by standard MD engines (e.g., OpenMM [13]) to execute MD simulations. This approach streamlines the parameterization workflow and enhances the transferability of the force field across diverse chemical environments.
Figure 1: Framework of ByteFF-Pol Given molecular graphs, the ByteFF GNN model predicts both bonded and non-bonded force field parameters. During the training stage, these parameters are combined with atomic coordinates in energy functions to compute decomposed energy terms, which are then fitted to reference labels from the ALMO-EDA calculations. Once trained, the force field parameters are obtained in a single inference step, and are used for simulations in standard MD softwares.
Similar to ByteFF [55], the GNN model is composed of three layers. First, a feature layer extracts information about atoms and bonds from molecular graphs to construct atom and bond embeddings ($x_{n}$ and $x_{e}$). Next, these embeddings are propagated through a multi-layer edge-augmented graph transformer (EGT)[24] to produce the hidden representations of atoms and bonds ($h_{n}$ and $h_{e}$), describing local chemical environments. Finally, a pooling layer processes these hidden representations to generate bonded and non-bonded force field parameters. The GNN model carefully considers molecular symmetries in its 2D topology, so the predicted force field parameters maintain these symmetries.
The ByteFF-Pol force field energy function, $U^{\mathrm{FF}}$, is partitioned into bonded and non-bonded components:
$U^{\mathrm{FF}}=U^{\mathrm{FF}}_{\mathrm{bonded}}+U^{\mathrm{FF}}_{\mathrm{non-bonded}}$ (1)
The bonded energy term, $U^{\mathrm{FF}}_{\mathrm{bonded}}$, is consistent with the functional forms used in ByteFF and GAFF2, encompassing standard bond, angle, proper dihedral, and improper dihedral potentials.
The non-bonded energy, $U^{\mathrm{FF}}_{\mathrm{non-bonded}}$, consists of five components: repulsion ($U^{\mathrm{FF}}_{\mathrm{rep}}$), dispersion ($U^{\mathrm{FF}}_{\mathrm{disp}}$), permanent electrostatic ($U^{\mathrm{FF}}_{\mathrm{est}}$), polarization ($U^{\mathrm{FF}}_{\mathrm{pol}}$), and charge transfer ($U^{\mathrm{FF}}_{\mathrm{ct}}$) terms:
$\begin{split}U^{\mathrm{FF}}_{\mathrm{non-bonded}}=&U^{\mathrm{FF}}_{\mathrm{rep}}(\bm{r};\epsilon^{\mathrm{rep}},\lambda^{\mathrm{rep}},r^{*})+U^{\mathrm{FF}}_{\mathrm{disp}}(\bm{r};C_{6},r^{*})+U^{\mathrm{FF}}_{\mathrm{est}}(\bm{r};q)\\ &+U^{\mathrm{FF}}_{\mathrm{pol}}(\bm{r};q,\alpha)+U^{\mathrm{FF}}_{\mathrm{ct}}(\bm{r};\epsilon^{\mathrm{ct}},\lambda^{\mathrm{ct}},r^{*}).\end{split}$ (2)
Here, $\bm{r}$ is the atomic coordinates and $r^{*}$, $\epsilon^{\mathrm{rep}}$, $\lambda^{\mathrm{rep}}$, $C_{6}$, $q$, $\alpha$, $\epsilon^{\mathrm{ct}}$ and $\lambda^{\mathrm{ct}}$ are force field parameters predicted by the GNN model. The specific formulations of these terms are provided in Section 3.2. Crucially, this non-bonded decomposition is designed to align with the energy components given by the ALMO-EDA method. This alignment allows the force field parameters to be trained by fitting the decomposed energy terms to the corresponding references derived from the ALMO-EDA calculations.
1.2 Fitting to ALMO-EDA labels
[…]
During training, ByteFF-Pol predicts the decomposed interaction energies of molecular dimers on the training set. These predictions are then fitted to the corresponding EDA labels to optimize the GNN model parameters. After training, the performance of ByteFF-Pol is verified on the validation set (Figure2 (a)). Despite the fixed functional form, ByteFF-Pol’s predictions show reasonable agreement with the DFT labels. The discrepancies in the lower interaction energy regions are minor, whereas those in higher interaction energy (i.e., highly repulsive) regions are comparatively large. However, given that the probability of sampling the high-energy conformations is extremely low under the Boltzmann distribution at ambient conditions, these larger deviations are expected to have a negligible impact on MD simulations.
To illustrate the fitting performance for the decomposed energy terms, two dimer scan examples are presented in Figure 2(b) and (c). The first example shows the interaction energy between two neutral molecules as a function of separation distance. In this system, the interactions are relatively weak and dominated by dispersion. Using a damped $C_{6}$ form, ByteFF-Pol accurately reproduces the dispersion interaction. The second example illustrates the interaction between a neutral molecule and a lithium ion, where electrostatic and polarization effects play central roles. ByteFF-Pol exhibits a slight overestimation of polarization interactions at short ranges; however, due to a cumulative cancellation of errors among the different terms, it still results in an accurate representation of the total interaction energy.
Figure 2: ByteFF-Pol is validated against DFT references. (a) Comparison of interaction energies predicted by ByteFF-Pol versus DFT on the validation set (a random sampled subset of 76 k conformations). (b, c) Dimer scan of (b) two EMC molecules and (c) an EMC molecule with a lithium ion (solid lines for ByteFF-Pol and dashed lines for DFT). (d) Cluster of a lithium ion surrounded by three DMC molecules and a FSI- anion. The conformations were first relaxed by DFT (light blue carbon) and then by ByteFF-Pol (grey carbon). (e, f) Decomposed binding energies of clusters with neutral molecules only (e) and clusters with neutral molecules and ions (f). The energy terms are accumulated so that the final point equals to the overall binding energy (solid lines for ByteFF-Pol and dashed lines for DFT). The corresponding components and conformations of the clusters are provided in Supplementary Section 2. The full names and SMILES strings corresponding to the molecular abbreviations are listed in Supplementary Table 1.
While ByteFF-Pol is trained exclusively on dimer data, its physically-motivated functional forms and decomposed training strategy make it transferable to many-body clusters. This transferability is demonstrated in Figure 2(d), which shows a lithium ion and its first solvation shell extracted from an MD simulation. The cluster’s geometries were first optimized with DFT at the B3LYP-D3(BJ)/def2-SVPD level [44], represented by the light blue carbon atoms. The conformations were then re-optimized using ByteFF-Pol, represented by the grey carbon atoms. During the force field optimization, the atomic positions were restrained with a weak harmonic force (1 kcal/mol/Å2) to ensure the system remained in the same local minimum. The resulting structures from both methods show excellent alignment, demonstrating the capability of ByteFF-Pol to accurately predict equilibrium geometries for complex clusters.
We then investigate the binding energy ($U_{\mathrm{bind}}$) of many-body clusters to validate ByteFF-Pol, a more comprehensive metric for molecular interactions. Unlike interaction energy, which is calculated for a fixed geometry, binding energy accounts for structural relaxation. It’s defined as the energy difference between the optimized cluster and its constituent molecules optimized in isolation:
$\begin{split}U_{\mathrm{bind}}^{\mathrm{DFT/FF}}&=U^{\mathrm{DFT/FF}}\left(\bm{r}^{\mathrm{DFT/FF}}_{\mathrm{cluster}}\right)-\sum_{i\in n_{\mathrm{mol}}}U^{\mathrm{DFT/FF}}\left(\bm{r}^{\mathrm{DFT/FF}}_{i,\mathrm{sep}}\right)\\ &=\left(U^{\mathrm{DFT/FF}}\left(\bm{r}^{\mathrm{DFT/FF}}_{\mathrm{cluster}}\right)-\sum_{i\in n_{\mathrm{mol}}}U^{\mathrm{DFT/FF}}\left(\bm{r}^{\mathrm{DFT/FF}}_{i,\mathrm{cluster}}\right)\right)\\ &+\left(\sum_{i\in n_{\mathrm{mol}}}U^{\mathrm{DFT/FF}}\left(\bm{r}^{\mathrm{DFT/FF}}_{i,\mathrm{cluster}}\right)-\sum_{i\in n_{\mathrm{mol}}}U^{\mathrm{DFT/FF}}\left(\bm{r}^{\mathrm{DFT/FF}}_{i,\mathrm{sep}}\right)\right)\\ &=U_{\mathrm{int}}^{\mathrm{DFT/FF}}+U_{\mathrm{conf}}^{\mathrm{DFT/FF}},\end{split}$ (3)
where $\bm{r}^{\mathrm{DFT/FF}}_{\mathrm{cluster}}$ is the atomic coordinates of the cluster optimized at the DFT or FF level, and $\bm{r}^{\mathrm{DFT/FF}}_{i,\mathrm{sep}}$ is the optimized coordinates of an isolated molecule $i$. The binding energy is thus a sum of the interaction energy ($U_{\mathrm{int}}$) and the conformational energy ($U_{\mathrm{conf}}$), which is the energy penalty associated with the conformation distortion of molecules going from isolated status into the cluster.
The binding energy of various clusters is shown in Figure 2 (e) for neutral systems and Figure 2 (f) for clusters containing ions. In these plots, the individual energy components are cumulatively summed so that the final value represents the total binding energy. The neutral examples were chosen to encompass diverse interaction types, including hydrogen bonding, polar interactions, $\pi-\pi$ stacking, etc. Similar to the dimer scan results, these interactions are relatively weak and dispersion-dominated. ByteFF-Pol exhibits a consistent trend with DFT calculations but tends to underestimate polarization and charge transfer energies, possibly due to the complex anisotropic nature of intermolecular interactions.
In clusters containing both neutral molecules and ions, the interactions are significantly stronger. Here, ByteFF-Pol demonstrates excellent agreement with DFT references. Notably, polarization interactions are inherently many-body effects, arising from the non-pairwise additive nature of electrical responses. Leveraging a physically-motivated self-consistent model for polarization, ByteFF-Pol maintains high accuracy in predicting this component even in extended clusters. Overall, the predictions from ByteFF-Pol for interaction energies, spanning from simple dimers to complex multi-molecule clusters, show strong concordance with the DFT benchmarks.
1.3 Thermodynamic Properties of Molecular Liquids
The thermodynamic properties of molecular liquids, primarily governed by intermolecular interactions, serve as critical benchmarks for evaluating force field validity. Using ByteFF-Pol, we computed the densities and evaporation enthalpies for a diverse set of nearly one hundred pure molecular liquids, covering a wide range of functional groups to represent different interaction types. The performance of ByteFF-Pol is compared against results from OPLS-AA [26, 25], OpenFF (i.g. OpenFF-unconstrained-2.2.1) [3], and MACE-OFF (i.g., MACE-OFF23(M)) [32], as shown in Figure 3(a) and (b).
For liquid densities, ByteFF-Pol provides accurate predictions comparable to established force fields like OPLS-AA and OpenFF. The minor disagreement with experimental values is mainly due to a slight, systematic overestimation. This overestimation is consistent with the expected impact of neglecting the nuclear quantum effects (NQE), which typically decreases the density of hydrogen-rich molecules by about 2% [33]. We address the NQE by performing Path-Integral Molecular Dynamics (PIMD) simulations [49], with detailed in Supplementary Section 8.6. As shown in Figure 3(a), ByteFF-Pol-NQE has effectively corrected the systematic overestimation of ByteFF-Pol, resulting lower mean absolute error (MAE) and mean absolute percentage error (MAPE) than other force fields.
Predicting the evaporation enthalpy ($\Delta H_{\text{vap}}$), however, is more challenging due to experimental uncertainties and the ideal gas approximation used for the gas phase simulations (see Supplementary Section 8.2). Despite these difficulties, ByteFF-Pol demonstrates remarkable accuracy, outperforming both OPLS-AA and OpenFF. With the inclusion of NQE, ByteFF-Pol-NQE yields even better results. This high accuracy is particularly noteworthy given that ByteFF-Pol is trained exclusively on DFT dimer data, whereas OPLS-AA and OpenFF are parameterized using experimental data including densities and evaporation enthalpies.
When compared to MACE-OFF, a SOTA ML force field trained purely on DFT data, ByteFF-Pol exhibits substantially lower prediction errors for both density and evaporation enthalpy. This discrepancy highlights the effectiveness of the physically-motivated model and training strategy employed by ByteFF-Pol. Additionally, due to its relatively simple functional form, ByteFF-Pol achieves a simulation speed of 50 ns/day for a 10,000-atom system on an L20 GPU, which is far more efficient than typical ML force fields (several ns/day) [32, 17].
Figure 3: Benchmarks on the thermodynamic properties of molecular liquids. Comparison of different force fields on (a) density and (b) evaporation enthalpy of pure molecular liquids from QUBEKit [23]; (c) density of binary molecular liquids from ThermoML [8]; and (d) density prediction by ByteFF-Pol on additional pure molecular liquids from CRC handbook [35]. The results of MACE-OFF were acquired from Ref 32. Pearson correlation coefficients (r) of each force field are provided in the legends.
Table 1: Statistics results of thermodynamic properties of molecular liquids.
Density (g/mL) from QUBEKit [23] $\Delta H_{\mathrm{vap}}$ (kcal/mol) from QUBEKit [23]
$n_{\mathrm{system}}$ MAE MAPE Pearson $n_{\mathrm{system}}$ MAE MAPE Pearson
OPLS-AA 96 0.024 2.5% 0.99 96 1.64 16.0% 0.85
OpenFF 96 0.024 2.3% 0.99 96 1.30 12.8% 0.87
MACE-OFF 93 0.086 9.4% 0.87 81 2.15 23.2% 0.86
ByteFF-Pol 96 0.029 3.0% 0.98 96 1.13 11.2% 0.90
ByteFF-Pol-NQE 96 0.023 2.3% 0.98 96 0.76 7.5% 0.92
Density (g/mL) from ThermoML [8] Density (g/mL) from CRC handbook [35]
$n_{\mathrm{system}}$ MAE MAPE Pearson $n_{\mathrm{system}}$ MAE MAPE Pearson
OPLS-AA 594 0.018 2.02% 0.98 - - - -
OpenFF 592 0.034 3.78% 0.95 - - - -
ByteFF-Pol 595 0.028 3.14% 0.99 2152 0.032 3.28% 0.98
The accuracy of ByteFF-Pol is further validated by its density predictions for binary molecular mixtures, as shown in Figure 3(c). In these systems, ByteFF-Pol demonstrates excellent accuracy, comparable to that of OPLS-AA and OpenFF. A slight, systematic overestimation of density across most mixtures can again be attributed to the neglect of NQE. However, due to the substantial computational cost of PIMD simulations, we did not perform PIMD simulations for this or subsequent benchmarks. Minor underestimations are observed in some high-density systems containing polyhalogenated molecules, which is likely due to the inherent limitations of the model’s functional form in capturing anisotropic interactions such as halogen bonds and $\sigma$-holes. Nevertheless, ByteFF-Pol demonstrates reliable overall performance, effectively modeling intermolecular forces across diverse chemical environments.
The transferability and universality of ByteFF-Pol are evaluated using density data for pure molecular liquids sourced from CRC handbook (Figure 3 (d)). Despite being trained on fewer than 400 molecules, ByteFF-Pol delivers precise density predictions for over 2,000 distinct molecules. The data points are colored based on the similarity of the local atomic chemical environments between the tested molecules and the training set, gauged by the similarity of the embedding vectors. As shown in Figure 3 (d), the densities of most molecules with large similarity to the training set can be predicted accurately by ByteFF-Pol. This observation underscores that ByteFF-Pol’s transferability stems from its ability to capture the chemical environments of atoms, rather than explicit training on entire molecular structures. This demonstrates that comprehensive coverage of atomic chemical environments in the training set is sufficient to enable accurate predictions across chemically diverse molecules. Overall, ByteFF-Pol demonstrates exceptional performance in predicting the thermodynamic properties of molecular liquids with promising transferability.
1.4 Applications in Liquid Electrolytes
One important application of organic liquid simulation is the design of electrolytes for batteries. In addition to thermodynamic properties, transport properties such as viscosity, conductivity, and self-diffusion coefficient are also critical for electrolytes. For typical electrolyte systems, ByteFF-Pol is compared with two conventional force fields, OPLS-AA and OPLS-AA-SC. Instead of the original OPLS-AA charges, OPLS-AA-SC employs CHELPG charges [6] scaled by a factor of 0.8, an empirical charge model widely used to approximate polarization in electrolyte systems.
Figure 4: Evaluation of ByteFF-Pol for predicting electrolyte properties. (a-c) Comparison of different force fields in predicting (a) density, (b) viscosity and (c) conductivity of liquid electrolytes, respectively. Pearson correlation coefficients (r) of each force field are provided in the legends. (d-e) Examples for conductivity at (d) varying lithium salt concentrations; (e) at different temperatures (solid lines denote simulation results, dashed lines denote experimental data). The numbers in the legends indicate the approximate molar fraction of each species. (f) Examples for self-diffusion coefficient ($D$) of different species in electrolyte. Upper bars correspond to $D$ of species X (labeled on the x axis) in the X:EMC binary solvent (molar ratio=1:1); and lower bars show the $D$ of X in the LiPF6:X:EMC electrolytes (molar ratio=1:4:4). Experimental data are obtained from [45]. The full names and SMILES strings corresponding to the molecular abbreviations are provided in Supplementary Table 1.
Table 2: Statistics results of electrolyte systems.
Density (g/mL) Viscosity (cP)
$n_{\mathrm{system}}$ MAE MAPE Pearson $n_{\mathrm{system}}$ MAE MAPE Pearson
OPLS-AA 88 0.022 1.76% 0.98 91 12.20 163.3% 0.77
OPLS-AA-SC 88 0.054 4.38% 0.97 91 1.79 39.8% 0.83
ByteFF-Pol 88 0.024 1.94% 0.99 91 1.06 20.0% 0.96
Conductivity (mS/cm) Conductivity >> 2.0 mS/cm
$n_{\mathrm{system}}$ MAE MAPE Pearson $n_{\mathrm{system}}$ MAE MAPE Pearson
OPLS-AA 4896 4.46 86.3% 0.66 3360 6.20 86.9% 0.57
OPLS-AA-SC 4896 1.55 64.9% 0.77 3360 2.04 29.5% 0.63
ByteFF-Pol 4896 0.86 52.4% 0.95 3360 1.09 17.1% 0.92
As shown in Figure 4 (a, b, and c), ByteFF-Pol exhibits excellent accuracy in predicting the density, viscosity, and conductivity of electrolyte systems, outperforming other evaluated force fields. OPLS-AA maintains high accuracy for density but overestimates viscosity and underestimates conductivity. While the scaled charge approach employed by OPLS-AA-SC improves conductivity accuracy, it introduces underestimation of both density and viscosity. In contrast, ByteFF-Pol accurately predicts all three properties, emphasizing the importance of explicitly modeling polarization. On this extensive conductivity benchmark encompassing nearly 5,000 electrolyte systems with diverse components and temperature conditions, ByteFF-Pol achieves a Pearson correlation coefficient of 0.95, demonstrating its utility as a tool for screening high-conductivity electrolytes. For other electrolyte force fields, large-scale benchmarking of this scope was not feasible; however, comparative analyses provided in Supplementary Section 9 underscore ByteFF-Pol’s advantages.
To further demonstrate the utility of ByteFF-Pol in electrolyte design, we present representative examples highlighting its practical potential. In real-world research, a key objective is to identify the optimal lithium salt concentration that maximizes electrolyte conductivity—a goal achievable by leveraging ByteFF-Pol to predict conductivity across varying lithium salt concentrations. Examples encompass systems containing two types of lithium salts, paired with cyclic carbonates, linear carbonates, esters, ethers, and nitriles, with conductivity values spanning 0.1 to $35\text{\,}\mathrm{mS}\text{\,}{\mathrm{cm}}^{-1}$ are provided in Figure 4 (d). It is shown that ByteFF-Pol not only delivers accurate conductivity predictions for diverse systems but also precisely captures the positions of conductivity peaks.
[…]
The diffusion coefficient of solvent molecules in electrolytes is a key property linked to the solvation structure of lithium ions [45]. Given that the definition and derivation of solvation structures lack clear standardization, we opted to directly benchmark ByteFF-Pol’s performance by comparing its predicted diffusion coefficients against experimental measurements. Figure 4(f) presents these coefficients for various solvents in systems with and without lithium salts. For common esters and carbonates, ByteFF-Pol’s predictions exhibit excellent agreement with experimental results. For fluorinated solvents, ByteFF-Pol shows a systematic underestimation, which, as discussed previously, can be attributed to its limitations in capturing anisotropic interactions. Despite this, ByteFF-Pol correctly predicts the experimental trends even for these challenging systems. In comparison, OPLS-AA consistently underestimates the diffusion coefficient across all systems, while OPLS-AA-SC consistently overestimates it.
In summary, ByteFF-Pol exhibits excellent accuracy in predicting diverse electrolyte properties. This capability, on one hand, validates the feasibility of developing universal force fields from ab initio calculations and, on the other hand, paves the way for exploring novel electrolytes across a vast chemical space.
2 Discussion
A longstanding goal in computational chemistry and materials science is to predict the macroscopic properties of condensed matter systems from ab initio computations. However, obtaining properties such as conductivity via ab initio MD simulations remains computationally prohibitive.
In this study, ByteFF-Pol serves as a critical bridge between microscopic DFT calculations and macroscopic liquid properties. Trained exclusively on DFT dimer labels, ByteFF-Pol demonstrates robust accuracy in predicting both thermodynamic properties (e.g., density and evaporation enthalpy) and transport properties (e.g., viscosity, diffusion coefficient, and conductivity), as validated through benchmarks covering extensive chemical space.
Its exceptional performance stems from two key foundations:
  1. 1.
    The physically-motivated design of its force field form and training scheme enables it to predict macroscopic properties despite being trained exclusively on QM dimer data.
  2. 2.
    The GNN-parametrization methodology, combined with efficient batch training on large-scale datasets, endows it with strong transferability across the chemical space of organic molecules, warranting its reliability in predicting properties of unseen molecules.
These advantages are pivotal to ByteFF-Pol’s success, rendering it a truly universal and ab initio force field that operates without reliance on experimental data. Its zero-shot predictive power and transferability open new avenues for research in functional liquid design, from electrolytes and deep eutectic solvents to other complex liquid systems.
While ByteFF-Pol already outperforms existing force fields, it retains potential for further optimization. Future work will focus on enhancing its functional form to better model anisotropic interactions, thereby improving accuracy in systems containing complex local interactions. In summary, ByteFF-Pol represents a significant step toward ab initio-guided prediction of macroscopic properties, offering a scalable, data-efficient tool for materials discovery.
3 Method
3.1 ALMO-EDA
The total interaction energy within a multi-fragment system, $U_{\mathrm{int}}^{\mathrm{DFT}}$, is calculated at the DFT level as the difference between the self-consistent field (SCF) energy of the entire cluster ($U^{\mathrm{DFT}}[\rho_{\mathrm{full}}]$) and the sum of the SCF energy of the isolated fragments ($U_{F}^{\mathrm{DFT}}[\rho_{F}]$):
$U_{\mathrm{int}}^{\mathrm{DFT}}=U^{\mathrm{DFT}}[\rho_{\mathrm{full}}]-\sum_{F}U_{F}^{\mathrm{DFT}}[\rho_{F}].$ (4)
This interaction energy is partitioned into physically meaningful components using the second-generation ALMO-EDA [28]. This variational scheme decomposes the total interaction energy as follows:
$U_{\mathrm{int}}^{\mathrm{DFT}}=U_{\mathrm{frz}}^{\mathrm{EDA}}+U_{\mathrm{pol}}^{\mathrm{EDA}}+U_{\mathrm{ct}}^{\mathrm{EDA}}.$ (5)
Here, $U_{\mathrm{frz}}^{\mathrm{EDA}}$ represents the frozen energy, which arises from the interaction between fragments without relaxing electronic structure. The polarization energy, $U_{\mathrm{pol}}^{\mathrm{EDA}}$, accounts for the energy stabilization as each fragment’s electron density relaxes in the electric field of its neighbors. Finally, the charge-transfer energy, $U_{\mathrm{ct}}^{\mathrm{EDA}}$, corresponds to the energy stabilization from electron delocalization between fragments.
The frozen term is further decomposed into dispersion ($U_{\mathrm{disp}}^{\mathrm{EDA}}$), electrostatic ($U_{\mathrm{est}}^{\mathrm{EDA}}$), and Pauli ($U_{\mathrm{pauli}}^{\mathrm{EDA}}$) terms [21]:
$U_{\mathrm{frz}}^{\mathrm{EDA}}=U_{\mathrm{disp}}^{\mathrm{EDA}}+U_{\mathrm{est}}^{\mathrm{EDA}}+U_{\mathrm{pauli}}^{\mathrm{EDA}}.$ (6)
The precise mathematical definitions and implementation details of all decomposed energy terms are provided in Supplementary Section 5.
3.2 Force Field Formulation
As discussed in Section 1.1, the ByteFF-Pol potential energy function, $U^{\mathrm{FF}}$, consists of bonded terms and non-bonded terms,
$U^{\mathrm{FF}}$ $=U_{\mathrm{bonded}}^{\mathrm{FF}}+U_{\mathrm{non-bonded}}^{\mathrm{FF}}.$ (7)
The non-bonded energy, $U_{\mathrm{non-bonded}}^{\mathrm{FF}}$, comprises five terms:
$U_{\mathrm{non-bonded}}^{\mathrm{FF}}$ $=U_{\mathrm{rep}}^{\mathrm{FF}}+U_{\mathrm{disp}}^{\mathrm{FF}}+U_{\mathrm{est}}^{\mathrm{FF}}+U_{\mathrm{pol}}^{\mathrm{FF}}+U_{\mathrm{ct}}^{\mathrm{FF}}.$ (8)
Repulsion and dispersion interactions are described by a modified Buckingham (exp-6) potential [38, 5], which has shown superior agreement with DFT-computed intermolecular energies compared to the Lennard-Jones potential [42, 5]. The functional forms are:
$U_{\mathrm{rep}}^{\mathrm{FF}}(\bm{r};\epsilon^{\mathrm{rep}},\lambda^{\mathrm{rep}},r^{*})$ $=\sum_{i<j}\left(\epsilon^{\mathrm{rep}}_{ij}\exp\left(\lambda^{\mathrm{rep}}_{ij}(1-\frac{r_{ij}}{r^{*}_{ij}})\right)+\left(\frac{s_{\mathrm{rep}}}{r_{ij}}\right)^{12}\right)$ (9)
$U_{\mathrm{disp}}^{\mathrm{FF}}(\bm{r};C_{6},r^{*})$ $=-\sum_{i<j}\frac{C_{6,ij}}{s_{\mathrm{disp}}{r^{*}_{ij}}^{6}+r_{ij}^{6}}.$ (10)
Here, $r_{ij}$ is the interatomic distance between atoms $i$ and $j$, $\epsilon^{\mathrm{rep}}$, $\lambda^{\mathrm{rep}}$, $C_{6}$, and $r^{*}$ are atom-specific parameters predicted by the GNN model. An additional $1/r^{12}$ term in the repulsion potential, scaled by a small constant $s_{\mathrm{rep}}$ (1.5 $\cdot(\mathrm{kcal/mol})^{1/12}$), is included to prevent unphysical attraction at very short interatomic distances. The dispersion term is damped at short ranges to avoid singularities and the double-counting of correlation effects, analogous to the DFT-D3 treatment [18]. The dimensionless damping factor $s_{\mathrm{disp}}$ is set to 120. As demonstrated in Section 1.2, this damped dispersion function provides an excellent fit to the target ALMO-EDA data.
The electrostatic interaction are modeled following the AMOEBA polarizable framework [11]. We use atom-centered point charges for the permanent electrostatics and an induced dipole model for polarization:
$U_{\mathrm{est}}^{\mathrm{FF}}(\bm{r};q)$ $=\sum_{i<j}\frac{q_{i}q_{j}}{4\pi\epsilon_{0}r_{ij}}$ (11)
$U_{\mathrm{pol}}^{\mathrm{FF}}(\bm{r};q,\alpha)$ $=-\frac{1}{2}\sum_{i}\boldsymbol{\mu}_{i}^{\mathrm{ind}}\cdot\mathbf{E}_{i}.$ (12)
This simplification from a full multipole expansion to monopoles streamlines parameterization without a significant loss of accuracy. The induced dipole, $\boldsymbol{\mu}^{\mathrm{ind}}$, is determined self-consistently by solving the polarization equations in response to the electric field $\mathbf{E}$. For efficient training, $\boldsymbol{\mu}^{\mathrm{ind}}$ is computed via direct matrix inversion, as detailed in Supplementary Section 6. The atomic charge $q$ and polarizability alphaalpha used in solving $\mu^{\mathrm{ind}}$ are predicted by the GNN model.
Finally, the charge-transfer term is modeled using a functional form inspired by the work of Deng et al. [12]:
$U_{\mathrm{ct}}^{\mathrm{FF}}(\bm{r};\epsilon^{\mathrm{ct}},\lambda^{\mathrm{ct}},r^{*})=\sum_{i<j}\frac{\epsilon^{\mathrm{ct}}_{ij}}{r_{ij}^{4}}\exp\left(-\left(\frac{\lambda^{\mathrm{ct}}_{ij}r^{*}_{ij}}{r_{ij}}\right)^{3}\right),$ (13)
The GNN model predicts the charge transfer parameters $\epsilon^{\mathrm{ct}}$ and $\lambda^{\mathrm{ct}}$. This pairwise additive functional form provides a good fit to the target ALMO-EDA data.
The atomic non-bonded parameters are predicted by the GNN model through several distinct approaches. Similar to our previous work on ByteFF, atomic charges $q$ are predicted using a Bond Charge Correction (BCC) model to preserve charge conservation. Furthermore, a subset of parameters are obtained using a physically-inspired Tkatchenko-Scheffler (TS) scaling model [48]:
$C_{6,i}=C_{6,i}^{\mathrm{free}}\left(\frac{V_{i,\mathrm{eff}}}{V_{i,\mathrm{free}}}\right)^{2},\ \alpha_{i}=\alpha_{i}^{\mathrm{free}}\left(\frac{V_{i,\mathrm{eff}}}{V_{i,\mathrm{free}}}\right),\ r_{i}^{*}=r_{i}^{*,\mathrm{free}}\left(\frac{V_{i,\mathrm{eff}}}{V_{i,\mathrm{free}}}\right)^{4/21}.$ (14)
In this model, $V_{i,\mathrm{eff}}$ is the effective atomic volume, predicted from the GNN’s atomic hidden states via a Multi-Layer Perceptron (MLP) to account for the local chemical environment. The terms $C_{6,i}^{\mathrm{free}}$, $\alpha_{i}^{\mathrm{free}}$, $r_{i}^{*,\mathrm{free}}$, and $V_{i,\mathrm{free}}$ are parameters for the free atom, dependent only on its element type. Finally, the remaining non-bonded parameters ($\epsilon^{\mathrm{rep}}$, $\lambda^{\mathrm{rep}}$, $\epsilon^{\mathrm{ct}}$, and $\lambda^{\mathrm{ct}}$) are predicted directly by MLPs that take the atom hidden states as input. We found that incorporating this physically-grounded TS model enhances training stability and model transferability without a noticeable loss of accuracy.
Once all atom-wise parameters are determined, the pairwise interaction parameters are derived using combination rules. We employ the standard Lorentz-Berthelot (LB) rules [36], which use an arithmetic mean for $r_{i}^{*}$ and a geometric mean for all other parameters:
$r_{ij}^{*}=\frac{r_{i}^{*}+r_{j}^{*}}{2},\quad\epsilon_{ij}=\sqrt{\epsilon_{i}\epsilon_{j}},\quad\lambda_{ij}=\sqrt{\lambda_{i}\lambda_{j}},\quad C_{6,ij}=\sqrt{C_{6,i}C_{6,j}}.$ (15)
3.3 Dataset
ByteFF-Pol’s training relies on three core datasets: the existing optimization and torsion datasets from our previous work and a newly created dimer dataset. Details on the original optimization and torsion datasets are in our previous publication [55]. For this work, we’ve augmented the optimization dataset by calculating the minimum basis iterative stockholder (MBIS) charges and volumes for each molecule at the PBE0/def2-TZVPD level, which are used to pre-train the model.
[…]
To benchmark the performance of ByteFF-Pol, we used five datasets of experimental physical properties. The first contains 96 pure molecular liquids with known densities and evaporation enthalpies from the QUBEKit benchmark [23]. The second and third datasets consist of experimental densities for 595 binary liquids from the NIST ThermoML archive [8] and 2132 pure molecular liquids from the CRC Handbook [35], respectively. For electrolyte-relevant systems, we compiled two additional datasets collected from literatures [54]: one with 88 experimental densities and 91 experimental viscosities for pure solvents and lithium electrolytes, and another containing 4896 conductivity measurements for various lithium electrolytes across different compositions and temperatures. The number of results for each force field may differ from that of the benchmark set, owing to parametrization errors or MD simulation failures.
3.4 Training Techniques
Our training procedure consists of three stages: pre-training, main training, and fine-tuning. Detailed training techniques are provided in Supplementary Section 7; here, we present a brief overview.
In the initial pre-training stage, the GNN model is trained to reproduce the bonded parameters from ByteFF [55], along with atomic charges and volumes from MBIS calculations. This stage allows the model to develop atomic hidden states that recognize the chemical environment and provides a physically-grounded starting point for the subsequent training.
In the main training stage, the 6 million data points are split into 5.4 million for training and 0.6 million for validation. We freeze most of the model and train only the output layers corresponding to the non-bonded interaction parameters. The training targets are the total interaction energies and their decomposed components from the ALMO-EDA calculations. Our force field’s energy decomposition scheme is carefully designed to align with the ALMO-EDA scheme, ensuring that the model learns physically meaningful interactions, as detailed in Supplementary Section 7.
Finally, in the fine-tuning stage, we update the torsion parameters to ensure they are consistent with the newly trained non-bonded interactions. The training target for this stage is the set of torsion scan energy profiles from ByteFF torsion dataset [55].
4 Data availability
The source data for Figure 3 and Figure 4 are provided in the supplementary information. A subset of the training dataset used in this study will be released with the source code.
5 Code availability
The code for ByteFF-Pol including the training scheme and trained model will be released upon acceptance of this paper.

### v4

@@
 Tianze Zheng at , Wen Yan at
-Molecular dynamics (MD) simulations have emerged as one of the most important cornerstones of modern materials and biological research, offering an indispensable tool for probing condensed-phase systems at the atomic level [29, 20, 53].
-By revealing microscopic details of complex phenomena such as structural rearrangements and molecular diffusion—often at resolutions inaccessible to experiments—MD simulations provide critical insights for applications ranging from drug discovery to material engineering [41, 47, 1].
-Among these applications, MD simulations play a vital role in the screening and design of organic molecules, such as electrolytes for novel batteries [16, 17] and deep eutectic solvents for green chemistry [50].
+Molecular dynamics (MD) simulations have emerged as one of the most important cornerstones of modern materials and biological research, offering an indispensable tool for probing condensed-phase systems at the atomic level [kholmurodov2007molecular, hollingsworth2018molecular, yan2023molecular].
+By revealing microscopic details of complex phenomena such as structural rearrangements and molecular diffusion—often at resolutions inaccessible to experiments—MD simulations provide critical insights for applications ranging from drug discovery to material engineering [reymond2012exploring, tafrishi2022molecular, alonso2006combining].
+Among these applications, MD simulations play a vital role in the screening and design of organic molecules, such as electrolytes for novel batteries [golovizninaExtensionCLPolPolarizable2021, gong2025predictive] and deep eutectic solvents for green chemistry [https://doi.org/10.1002/wcms.1598].
 They enable the rapid exploration of vast chemical and formula spaces for organic materials, which is impossible to enumerate through experiments.
@@
 This dependency necessitates the development of universal force fields combining accuracy and efficiency.
-Traditional force fields, such as Amber [22], CHARMM [7], and OPLS [26, 25], model interatomic interactions using a set of simple and predefined functional forms and tabulated parameters.
+Traditional force fields, such as Amber [hornak2006comparison], CHARMM [brooks2009charmm], and OPLS [jorgensen1984optimized, jorgensen1986optimized], model interatomic interactions using a set of simple and predefined functional forms and tabulated parameters.
 Conventional wisdom indicates that the accuracy of traditional force fields for macroscopic properties essentially relies on error cancellation at microscopic level, a consequence of their limited expressive power.
-To achieve this cancellation, the force field parameters derived from low-level quantum mechanics (QM) calculations need to be further optimized by experimental data, including spectroscopic measurements [19], and thermodynamic properties such as densities and evaporation enthalpies [26, 25].
-In contrast, machine learning (ML)-based force fields utilize the powerful fitting capabilities of neural networks to directly learn interatomic interactions from QM calculations, thereby achieving high accuracy in predicting intramolecular energy landscapes, spectroscopic properties, and related phenomena [31, 32, 27, 39].
+To achieve this cancellation, the force field parameters derived from low-level quantum mechanics (QM) calculations need to be further optimized by experimental data, including spectroscopic measurements [han2025refining], and thermodynamic properties such as densities and evaporation enthalpies [jorgensen1984optimized, jorgensen1986optimized].
+In contrast, machine learning (ML)-based force fields utilize the powerful fitting capabilities of neural networks to directly learn interatomic interactions from QM calculations, thereby achieving high accuracy in predicting intramolecular energy landscapes, spectroscopic properties, and related phenomena [kovacsEvaluationMACEForce2023, kovacsMaceOff2025, juSevenNet2025, pleFoundationModelAccurate2025].
 However, comparing to traditional force fields, ML force fields without physical constraints require tremendous QM data for training.
 The development of a universal many-body ML force field is exceptionally challenging due to the vast chemical space of organic molecules and the daunting computational cost of generating sufficient ab initio data.
-Consequently, prominent ML force fields like MACE-OFF can suffer from accuracy and transferability issues, resulting in inferior performance in bulk property predictions when compared to well-established traditional ones [32].
-While other efforts, such as the BAMBOO force field  [17], have demonstrated state-of-the-art (SOTA) performance in predicting the bulk properties of electrolytes, they still depend on experimental density data for empirical fine-tuning [17].
+Consequently, prominent ML force fields like MACE-OFF can suffer from accuracy and transferability issues, resulting in inferior performance in bulk property predictions when compared to well-established traditional ones [kovacsMaceOff2025].
+While other efforts, such as the BAMBOO force field [gong2025predictive], have demonstrated state-of-the-art (SOTA) performance in predicting the bulk properties of electrolytes, they still depend on experimental density data for empirical fine-tuning [gong2025predictive].
 Ultimately, the predictive capability of both traditional and ML force fields remains constrained by the availability and integration of experimental information.
-To address this challenge, efforts have been made toward establishing ab initio force fields by refining traditional functional forms to better incorporate the underlying physics of molecular interactions [52, 23].
-These force fields are typically based on separating the intermolecular interactions into distinct components [43, liu2019amoeba+] or employing many-body expansion [2].
-Despite the solid physical foundation, such force fields often struggle with computational efficiency and transferability [52].
-Polarizable force fields, including AMOEBA [40, 9, liu2019amoeba+], APPLE&P [4] and CL&Pol [16], have emerged as a successful framework that retains computational efficiency while capturing the response of electron density to different electrical environments—a critical feature for systems like electrolytes [30].
+To address this challenge, efforts have been made toward establishing ab initio force fields by refining traditional functional forms to better incorporate the underlying physics of molecular interactions [xu2018perspective, hortonQUBEKitAutomatingDerivation2019].
+These force fields are typically based on separating the intermolecular interactions into distinct components [schmidt_transferable_2015, liu2019amoeba+] or employing many-body expansion [beran2009approximating].
+Despite the solid physical foundation, such force fields often struggle with computational efficiency and transferability [xu2018perspective].
+Polarizable force fields, including AMOEBA [ponderCurrentStatusAMOEBA2010, chungClassicalExchangePolarization2022, liu2019amoeba+], APPLE&P [borodinPolarizableForceField2009] and CL&Pol [golovizninaExtensionCLPolPolarizable2021], have emerged as a successful framework that retains computational efficiency while capturing the response of electron density to different electrical environments—a critical feature for systems like electrolytes [kirshCriticalEvaluationPolarizable2024].
 Nevertheless, parameterizing these polarizable force fields remains a complex, labor-intensive process that often relies on manual adjustments.
 As a result, their transferability across different chemical environments is uncertain, failing to consistently outperform traditional force fields when applied to systems with varying electronic structures or bonding characteristics.
-In summary, currently there is still an urgent need for an universal organic force field that possesses both the accuracy of high-level ab initio calculations, as well as running speed, generalizability, and data efficiency required by large-scale simulations of complex materials.
-In response to this challenge, we introduce ByteFF-Pol, a graph neural network [10] (GNN)-parameterized polarizable force field that represents a significant advancement in force field development.
+In summary, currently there is still an urgent need for a universal organic force field that possesses both the accuracy of high-level ab initio calculations, as well as running speed and generalizability required by large-scale simulations of complex materials.
+In response to this challenge, we introduce ByteFF-Pol, a graph neural network [corso2024graph] (GNN)-parameterized polarizable force field that represents a significant advancement in force field development.
 Unlike traditional and most ML-based force fields, ByteFF-Pol is trained exclusively on high-level QM labels, eliminating the need for experimental calibration.
-Through a novel force field architecture and training scheme, ByteFF-Pol achieves exceptional accuracy in predicting the thermodynamic and transport properties of small-molecule liquids and electrolytes, outperforming SOTA traditonal and ML force fields.
+Through a novel force field architecture and training scheme, ByteFF-Pol achieves exceptional accuracy in predicting the thermodynamic and transport properties of small-molecule liquids and electrolytes, outperforming SOTA traditional and ML force fields.
 Its zero-shot prediction capabilities enable the exploration of previously intractable chemical spaces, opening up new avenues for data-driven discovery of materials with optimized properties for applications in electrolyte design, complex solvents, and beyond.
@@
 1.1 Overview of ByteFF-Pol
-The framework of ByteFF-Pol is illustrated in Figure 1.
-Unlike traditional methods that rely on look-up tables, ByteFF-Pol employs a GNN model to predict force field parameters directly from molecular graphs.
-During the training stage, these parameters are integrated with atomic coordinates into energy functions to compute the decomposed energy terms, which are then fitted to reference labels from the absolutely localized molecular orbital energy decomposition analysis (ALMO-EDA) calculations [28].
-After training, the parameters of the GNN model are fixed.
-The force field parameters predicted by the GNN model can be directly used by standard MD engines (e.g., OpenMM [13]) to execute MD simulations.
+The ByteFF-Pol framework, depicted in Figure 1, utilizes a Graph Neural Network (GNN) to derive force field parameters directly from molecular graphs.
+This modern parameterization technique offers significant advantages over traditional look-up tables [wangEndtoendDifferentiableConstruction2022, thurlemann2023regularized, zheng2025data, seute2025grappa].
+During training, the predicted force field parameters are integrated with atomic coordinates into energy functions to compute the decomposed energy terms.
+And the GNN model is optimized by fitting the decomposed energy terms to reference data from the absolutely localized molecular orbital energy decomposition analysis (ALMO-EDA) calculations [khaliulinEDAv1Original2007].
+Once trained, the finalized GNN model can generate parameters for direct use in standard molecular dynamics (MD) engines (e.g., OpenMM [eastman2017openmm]) to execute MD simulations.
 This approach streamlines the parameterization workflow and enhances the transferability of the force field across diverse chemical environments.
@@
 Once trained, the force field parameters are obtained in a single inference step, and are used for simulations in standard MD softwares.
-Similar to ByteFF [55], the GNN model is composed of three layers.
+Similar to ByteFF [zheng2025data], the GNN model is composed of three layers.
 First, a feature layer extracts information about atoms and bonds from molecular graphs to construct atom and bond embeddings ($x_{n}$ and $x_{e}$).
-Next, these embeddings are propagated through a multi-layer edge-augmented graph transformer (EGT)[24] to produce the hidden representations of atoms and bonds ($h_{n}$ and $h_{e}$), describing local chemical environments.
+Next, these embeddings are propagated through a multi-layer edge-augmented graph transformer (EGT)[hussainGlobalSelfAttentionReplacement2022] to produce the hidden representations of atoms and bonds ($h_{n}$ and $h_{e}$), describing local chemical environments.
 Finally, a pooling layer processes these hidden representations to generate bonded and non-bonded force field parameters.
@@
 This alignment allows the force field parameters to be trained by fitting the decomposed energy terms to the corresponding references derived from the ALMO-EDA calculations.
+ByteFF-Pol shares a similar GNN model with our previous work, ByteFF, for predicting force field parameters.
+The major advancements lie in the parametrization methodology of non-bonded parameters: instead of distilling the conventional GAFF2 force field, we now fit to QM data directly.
+This constitutes a non-trivial task, as most SOTA force fields require experimental fine-tuning due to their inaccurate function form and incomplete QM dataset.
+To address these challenges, we curated a large-scale, chemically diverse QM dataset, generating decomposed energy labels using the ALMO-EDA method in addition to the total energy.
+The EDA data encodes physically meaningful information that substantially enhances the robustness of the resulting model.
+Correspondingly, we designed EDA-compatible force field functions and a tailored training protocol to ensure both fitting efficiency and accuracy.
+In the subsequent sections, we detail the training procedure of ByteFF-Pol and demonstrate its exceptional performance in both microscopic energy predictions and bulk simulations through comprehensive benchmarking studies.
 1.2 Fitting to ALMO-EDA labels
@@
 The energy terms are accumulated so that the final point equals to the overall binding energy (solid lines for ByteFF-Pol and dashed lines for DFT).
-The corresponding components and conformations of the clusters are provided in Supplementary Section 2.
+The corresponding components and conformations of the clusters are provided in Supplementary Section 2.1.
 The full names and SMILES strings corresponding to the molecular abbreviations are listed in Supplementary Table 1.
@@
 This transferability is demonstrated in Figure 2(d), which shows a lithium ion and its first solvation shell extracted from an MD simulation.
-The cluster’s geometries were first optimized with DFT at the B3LYP-D3(BJ)/def2-SVPD level [44], represented by the light blue carbon atoms.
+The cluster’s geometries were first optimized with DFT at the B3LYP-D3(BJ)/def2-SVPD level [stephens1994ab], represented by the light blue carbon atoms.
 The conformations were then re-optimized using ByteFF-Pol, represented by the grey carbon atoms.
@@
 Using ByteFF-Pol, we computed the densities and evaporation enthalpies for a diverse set of nearly one hundred pure molecular liquids, covering a wide range of functional groups to represent different interaction types.
-The performance of ByteFF-Pol is compared against results from OPLS-AA [26, 25], OpenFF (i.g.
-OpenFF-unconstrained-2.2.1) [3], and MACE-OFF (i.g., MACE-OFF23(M)) [32], as shown in Figure 3(a) and (b).
-For liquid densities, ByteFF-Pol provides accurate predictions comparable to established force fields like OPLS-AA and OpenFF.
+The performance of ByteFF-Pol is benchmarked against that of SOTA force fields, including a conventional force field (i.e.
+OPLS-AA [jorgensen1984optimized, jorgensen1986optimized]), a polarizable force field (i.e.
+AMOEBA [ponderCurrentStatusAMOEBA2010, walkerAutomationAMOEBAPolarizable2022]), and a ML force field (i.e., MACE-OFF23(M)) [kovacsMaceOff2025], as shown in Figure 3(a) and (b).
+Parmetrization and simulation details can be found in Supplementary Section 8, and comparsion with additional force fields are provided in Supplementary Section 9.1.
+For liquid densities, ByteFF-Pol provides accurate predictions comparable to established force fields like OPLS-AA and AMOEBA.
 The minor disagreement with experimental values is mainly due to a slight, systematic overestimation.
-This overestimation is consistent with the expected impact of neglecting the nuclear quantum effects (NQE), which typically decreases the density of hydrogen-rich molecules by about 2% [33].
-We address the NQE by performing Path-Integral Molecular Dynamics (PIMD) simulations [49], with detailed in Supplementary Section 8.6.
-As shown in Figure 3(a), ByteFF-Pol-NQE has effectively corrected the systematic overestimation of ByteFF-Pol, resulting lower mean absolute error (MAE) and mean absolute percentage error (MAPE) than other force fields.
+This overestimation is consistent with the expected impact of neglecting the nuclear quantum effects (NQEs), which typically decreases the density of hydrogen-rich molecules by about 2% [kurnikov2024neural].
+We address the NQEs by performing Path-Integral Molecular Dynamics (PIMD) simulations [tuckerman_statistical_2023], with detailed in Supplementary Section 8.6.
+As shown in Figure 3(a), ByteFF-Pol-NQE has effectively corrected the systematic overestimation of ByteFF-Pol, resulting a lower mean absolute error (MAE) and mean absolute percentage error (MAPE) than other force fields.
 Predicting the evaporation enthalpy ($\Delta H_{\text{vap}}$), however, is more challenging due to experimental uncertainties and the ideal gas approximation used for the gas phase simulations (see Supplementary Section 8.2).
-Despite these difficulties, ByteFF-Pol demonstrates remarkable accuracy, outperforming both OPLS-AA and OpenFF.
-With the inclusion of NQE, ByteFF-Pol-NQE yields even better results.
-This high accuracy is particularly noteworthy given that ByteFF-Pol is trained exclusively on DFT dimer data, whereas OPLS-AA and OpenFF are parameterized using experimental data including densities and evaporation enthalpies.
+Despite these difficulties, ByteFF-Pol demonstrates remarkable accuracy, outperforming both OPLS-AA and AMOEBA.
+With the inclusion of NQEs, ByteFF-Pol-NQE yields even better results.
+This high accuracy is particularly noteworthy given that ByteFF-Pol is trained exclusively on DFT dimer data, whereas OPLS-AA and AMOEBA are parameterized using experimental data including densities and evaporation enthalpies.
 When compared to MACE-OFF, a SOTA ML force field trained purely on DFT data, ByteFF-Pol exhibits substantially lower prediction errors for both density and evaporation enthalpy.
+As shown in Supplementary Figure 14, equipped with physics-driven long-range interactions, MACELES-OFF [kim2025universal] significantly outperforms MACE-OFF, although it still exhibits larger errors than ByteFF-Pol.
 This discrepancy highlights the effectiveness of the physically-motivated model and training strategy employed by ByteFF-Pol.
-Additionally, due to its relatively simple functional form, ByteFF-Pol achieves a simulation speed of 50 ns/day for a 10,000-atom system on an L20 GPU, which is far more efficient than typical ML force fields (several ns/day) [32, 17].
 Figure 3: Benchmarks on the thermodynamic properties of molecular liquids.
-Comparison of different force fields on (a) density and (b) evaporation enthalpy of pure molecular liquids from QUBEKit [23]; (c) density of binary molecular liquids from ThermoML [8]; and (d) density prediction by ByteFF-Pol on additional pure molecular liquids from CRC handbook [35].
-The results of MACE-OFF were acquired from Ref 32.
+Comparison of different force fields on (a) density and (b) evaporation enthalpy of pure molecular liquids from QUBEKit [hortonQUBEKitAutomatingDerivation2019]; (c) density of binary molecular liquids from ThermoML [chirico2003thermoml]; and (d) density prediction by ByteFF-Pol on additional pure molecular liquids from CRC handbook [Lide2004CRC].
+The results of MACE-OFF were acquired from Ref kovacsMaceOff2025.
 Pearson correlation coefficients (r) of each force field are provided in the legends.
 Table 1: Statistics results of thermodynamic properties of molecular liquids.
-Density (g/mL) from QUBEKit [23] $\Delta H_{\mathrm{vap}}$ (kcal/mol) from QUBEKit [23]
+Density (g/mL) from QUBEKit [hortonQUBEKitAutomatingDerivation2019] $\Delta H_{\mathrm{vap}}$ (kcal/mol) from QUBEKit [hortonQUBEKitAutomatingDerivation2019]
 $n_{\mathrm{system}}$ MAE MAPE Pearson $n_{\mathrm{system}}$ MAE MAPE Pearson
-OPLS-AA 96 0.024 2.5% 0.99 96 1.64 16.0% 0.85
-OpenFF 96 0.024 2.3% 0.99 96 1.30 12.8% 0.87
-MACE-OFF 93 0.086 9.4% 0.87 81 2.15 23.2% 0.86
-ByteFF-Pol 96 0.029 3.0% 0.98 96 1.13 11.2% 0.90
+OPLS-AA 96 0.024 2.5% 0.99 96 1.52 14.8% 0.87
+AMOEBA 87 0.025 2.4% 0.99 87 0.85 8.1% 0.91
+MACE-OFF 92 0.084 9.8% 0.92 80 2.14 23.1% 0.86
+ByteFF-Pol 96 0.029 3.0% 0.98 96 0.95 9.3% 0.94
 ByteFF-Pol-NQE 96 0.023 2.3% 0.98 96 0.76 7.5% 0.92
-Density (g/mL) from ThermoML [8] Density (g/mL) from CRC handbook [35]
+Density (g/mL) from ThermoML [chirico2003thermoml] Density (g/mL) from CRC handbook [Lide2004CRC]
 $n_{\mathrm{system}}$ MAE MAPE Pearson $n_{\mathrm{system}}$ MAE MAPE Pearson
-OPLS-AA 594 0.018 2.02% 0.98 - - - -
-OpenFF 592 0.034 3.78% 0.95 - - - -
-ByteFF-Pol 595 0.028 3.14% 0.99 2152 0.032 3.28% 0.98
+OPLS-AA 594 0.018 2.0% 0.98 - - - -
+AMOEBA 554 0.018 2.0% 0.99 - - - -
+ByteFF-Pol 595 0.028 3.1% 0.99 2152 0.032 3.28% 0.98
+ByteFF-Pol-NQE 595 0.017 1.8% 0.99 - - - -
 The accuracy of ByteFF-Pol is further validated by its density predictions for binary molecular mixtures, as shown in Figure 3(c).
-In these systems, ByteFF-Pol demonstrates excellent accuracy, comparable to that of OPLS-AA and OpenFF.
-A slight, systematic overestimation of density across most mixtures can again be attributed to the neglect of NQE.
+In these systems, ByteFF-Pol demonstrates excellent accuracy, comparable to that of OPLS-AA and AMOEBA.
+A slight, systematic overestimation of density by ByteFF-Pol can again be corrected by incorporating the NQEs.
 However, due to the substantial computational cost of PIMD simulations, we did not perform PIMD simulations for this or subsequent benchmarks.
@@
 As shown in Figure 3 (d), the densities of most molecules with large similarity to the training set can be predicted accurately by ByteFF-Pol.
+As demonstrated in Supplementary Section 3.2, a clear correlation exists between the similarity score and the density prediction accuracy.
 This observation underscores that ByteFF-Pol’s transferability stems from its ability to capture the chemical environments of atoms, rather than explicit training on entire molecular structures.
-This demonstrates that comprehensive coverage of atomic chemical environments in the training set is sufficient to enable accurate predictions across chemically diverse molecules.
+This indicates that comprehensive coverage of atomic chemical environments in the training set is sufficient to enable accurate predictions across chemically diverse molecules (see Supplementary Section 10 for a detailed analysis of the training set’s chemical space).
 Overall, ByteFF-Pol demonstrates exceptional performance in predicting the thermodynamic properties of molecular liquids with promising transferability.
@@
 In addition to thermodynamic properties, transport properties such as viscosity, conductivity, and self-diffusion coefficient are also critical for electrolytes.
-For typical electrolyte systems, ByteFF-Pol is compared with two conventional force fields, OPLS-AA and OPLS-AA-SC.
-Instead of the original OPLS-AA charges, OPLS-AA-SC employs CHELPG charges [6] scaled by a factor of 0.8, an empirical charge model widely used to approximate polarization in electrolyte systems.
+For typical electrolyte systems, ByteFF-Pol is compared with two conventional force fields, OPLS-AA and OPLS-AA-SC, as well as a SOTA polarizable force field, AMOEBA.
+Instead of the original OPLS-AA charges, OPLS-AA-SC employs CHELPG charges [breneman1990determining] scaled by a factor of 0.8, an empirical charge model widely used to approximate polarization in electrolyte systems.
 Figure 4: Evaluation of ByteFF-Pol for predicting electrolyte properties.
@@
 Upper bars correspond to $D$ of species X (labeled on the x axis) in the X:EMC binary solvent (molar ratio=1:1); and lower bars show the $D$ of X in the LiPF6:X:EMC electrolytes (molar ratio=1:4:4).
-Experimental data are obtained from [45].
+Experimental data are obtained from [C9EE00141G].
 The full names and SMILES strings corresponding to the molecular abbreviations are provided in Supplementary Table 1.
@@
 $n_{\mathrm{system}}$ MAE MAPE Pearson $n_{\mathrm{system}}$ MAE MAPE Pearson
-OPLS-AA 88 0.022 1.76% 0.98 91 12.20 163.3% 0.77
-OPLS-AA-SC 88 0.054 4.38% 0.97 91 1.79 39.8% 0.83
-ByteFF-Pol 88 0.024 1.94% 0.99 91 1.06 20.0% 0.96
+OPLS-AA 88 0.022 1.8% 0.98 91 12.20 163.3% 0.77
+OPLS-AA-SC 88 0.054 4.4% 0.97 91 1.79 39.8% 0.83
+AMOEBA 88 0.030 2.4% 0.98 90 6.69 103.4% 0.79
+ByteFF-Pol 88 0.024 1.9% 0.99 91 1.06 20.0% 0.96
 Conductivity (mS/cm) Conductivity >> 2.0 mS/cm
@@
 OPLS-AA-SC 4896 1.55 64.9% 0.77 3360 2.04 29.5% 0.63
+AMOEBA 3735 2.70 68.4% 0.93 2582 3.66 56.4% 0.90
 ByteFF-Pol 4896 0.86 52.4% 0.95 3360 1.09 17.1% 0.92
-As shown in Figure 4 (a, b, and c), ByteFF-Pol exhibits excellent accuracy in predicting the density, viscosity, and conductivity of electrolyte systems, outperforming other evaluated force fields.
+As shown in Figure 4, ByteFF-Pol exhibits excellent accuracy in predicting the density, viscosity, and conductivity of electrolyte systems, outperforming other evaluated force fields.
 OPLS-AA maintains high accuracy for density but overestimates viscosity and underestimates conductivity.
 While the scaled charge approach employed by OPLS-AA-SC improves conductivity accuracy, it introduces underestimation of both density and viscosity.
+These discrepancies primarily arise from the neglect of explicit polarization in conventional force fields, a factor demonstrated to be critical for electrolyte systems in the ablation study (Supplementary Section 11).
+Although AMOEBA achieves better alignment with experimental reference by accounting for polarization, it still exhibits a systematic overestimation of viscosity and underestimation of conductivity.
+This limitation might be attributed to its training protocol, which relies on fitting to experimental bulk data; consequently, the force field’s performance may diminish when the chemical space of the electrolyte systems extends beyond the coverage of its original experimental training set.
+Furthermore, AMOEBA suffers from occasional parameterization failures and simulation instabilities, resulting in unavailable conductivity data for approximately one-quarter of the tested systems.
 In contrast, ByteFF-Pol accurately predicts all three properties, emphasizing the importance of explicitly modeling polarization.
-On this extensive conductivity benchmark encompassing nearly 5,000 electrolyte systems with diverse components and temperature conditions, ByteFF-Pol achieves a Pearson correlation coefficient of 0.95, demonstrating its utility as a tool for screening high-conductivity electrolytes.
-For other electrolyte force fields, large-scale benchmarking of this scope was not feasible; however, comparative analyses provided in Supplementary Section 9 underscore ByteFF-Pol’s advantages.
+On this extensive conductivity benchmark encompassing nearly 5,000 electrolyte systems with diverse components and temperature conditions, ByteFF-Pol achieves a Pearson correlation coefficient of 0.95, while yielding the lowest MAE and MAPE, demonstrating its utility as a tool for screening high-conductivity electrolytes.
+For other electrolyte force fields, large-scale benchmarking of this scope was not feasible; however, comparative analyses provided in Supplementary Section 9.2 underscore ByteFF-Pol’s advantages.
+As shown in Supplementary Figure 15 and Supplementary Table 11, ByteFF-Pol exhibits superior accuracy in predicting conductivity compared to BAMBOO, the current SOTA machine learning force field for electrolyte systems.
+The experimental conductivity data presented in Supplementary Figure 15 were obtained from iterative optimization of an electrolyte system [zhu2024differentiable].
+As a result of the excellent agreement between ByteFF-Pol’s predictions and experimental observations, the model holds significant potential to replace such experimental measurements, thereby reducing associated costs.
 To further demonstrate the utility of ByteFF-Pol in electrolyte design, we present representative examples highlighting its practical potential.
 In real-world research, a key objective is to identify the optimal lithium salt concentration that maximizes electrolyte conductivity—a goal achievable by leveraging ByteFF-Pol to predict conductivity across varying lithium salt concentrations.
-Examples encompass systems containing two types of lithium salts, paired with cyclic carbonates, linear carbonates, esters, ethers, and nitriles, with conductivity values spanning 0.1 to $35\text{\,}\mathrm{mS}\text{\,}{\mathrm{cm}}^{-1}$ are provided in Figure 4 (d).
+Examples encompass systems containing two types of lithium salts, paired with cyclic carbonates, linear carbonates, esters, ethers, and nitriles, with conductivity values spanning 0.1 to 35 mS/cm are provided in Figure 4 (d).
 It is shown that ByteFF-Pol not only delivers accurate conductivity predictions for diverse systems but also precisely captures the positions of conductivity peaks.
 […]
-The diffusion coefficient of solvent molecules in electrolytes is a key property linked to the solvation structure of lithium ions [45].
+The diffusion coefficient of solvent molecules in electrolytes is a key property linked to the solvation structure of lithium ions [C9EE00141G].
 Given that the definition and derivation of solvation structures lack clear standardization, we opted to directly benchmark ByteFF-Pol’s performance by comparing its predicted diffusion coefficients against experimental measurements.
@@
 In comparison, OPLS-AA consistently underestimates the diffusion coefficient across all systems, while OPLS-AA-SC consistently overestimates it.
-In summary, ByteFF-Pol exhibits excellent accuracy in predicting diverse electrolyte properties.
+Owing to its relatively simple functional form, ByteFF-Pol achieves a simulation speed of approximately 40 ns/day for a 10,000-atom system on an A100 GPU, and 32 ns/day on the more cost-effective A10 GPU.
+This computational efficiency is markedly superior to that of typical ML force fields (see Supplementary Section 9.3 for detailed comparison).
+In summary, ByteFF-Pol exhibits excellent accuracy and efficiency in predicting diverse electrolyte properties.
 This capability, on one hand, validates the feasibility of developing universal force fields from ab initio calculations and, on the other hand, paves the way for exploring novel electrolytes across a vast chemical space.
@@
 $U_{\mathrm{int}}^{\mathrm{DFT}}=U^{\mathrm{DFT}}[\rho_{\mathrm{full}}]-\sum_{F}U_{F}^{\mathrm{DFT}}[\rho_{F}].$ (4)
-This interaction energy is partitioned into physically meaningful components using the second-generation ALMO-EDA [28]. This variational scheme decomposes the total interaction energy as follows:
+This interaction energy is partitioned into physically meaningful components using the second-generation ALMO-EDA [khaliulinEDAv1Original2007].
+This variational scheme decomposes the total interaction energy as follows:
 $U_{\mathrm{int}}^{\mathrm{DFT}}=U_{\mathrm{frz}}^{\mathrm{EDA}}+U_{\mathrm{pol}}^{\mathrm{EDA}}+U_{\mathrm{ct}}^{\mathrm{EDA}}.$ (5)
@@
 Finally, the charge-transfer energy, $U_{\mathrm{ct}}^{\mathrm{EDA}}$, corresponds to the energy stabilization from electron delocalization between fragments.
-The frozen term is further decomposed into dispersion ($U_{\mathrm{disp}}^{\mathrm{EDA}}$), electrostatic ($U_{\mathrm{est}}^{\mathrm{EDA}}$), and Pauli ($U_{\mathrm{pauli}}^{\mathrm{EDA}}$) terms [21]:
+The frozen term is further decomposed into dispersion ($U_{\mathrm{disp}}^{\mathrm{EDA}}$), electrostatic ($U_{\mathrm{est}}^{\mathrm{EDA}}$), and Pauli ($U_{\mathrm{pauli}}^{\mathrm{EDA}}$) terms [hornEDAElectrostaticDispersion2016]:
 $U_{\mathrm{frz}}^{\mathrm{EDA}}=U_{\mathrm{disp}}^{\mathrm{EDA}}+U_{\mathrm{est}}^{\mathrm{EDA}}+U_{\mathrm{pauli}}^{\mathrm{EDA}}.$ (6)
@@
 $U^{\mathrm{FF}}$ $=U_{\mathrm{bonded}}^{\mathrm{FF}}+U_{\mathrm{non-bonded}}^{\mathrm{FF}}.$ (7)
-The non-bonded energy, $U_{\mathrm{non-bonded}}^{\mathrm{FF}}$, comprises five terms:
+The detailed bonded terms could be found in Supplementary Section 6.2. The non-bonded energy, $U_{\mathrm{non-bonded}}^{\mathrm{FF}}$, comprises five terms:
 $U_{\mathrm{non-bonded}}^{\mathrm{FF}}$ $=U_{\mathrm{rep}}^{\mathrm{FF}}+U_{\mathrm{disp}}^{\mathrm{FF}}+U_{\mathrm{est}}^{\mathrm{FF}}+U_{\mathrm{pol}}^{\mathrm{FF}}+U_{\mathrm{ct}}^{\mathrm{FF}}.$ (8)
-Repulsion and dispersion interactions are described by a modified Buckingham (exp-6) potential [38, 5], which has shown superior agreement with DFT-computed intermolecular energies compared to the Lennard-Jones potential [42, 5].
+Repulsion and dispersion interactions are described by a modified Buckingham (exp-6) potential [mason_transport_1954, borodin_development_2006], which has shown superior agreement with DFT-computed intermolecular energies compared to the Lennard-Jones potential [rowley_determination_2001, borodin_development_2006].
 The functional forms are:
@@
 An additional $1/r^{12}$ term in the repulsion potential, scaled by a small constant $s_{\mathrm{rep}}$ (1.5 $\cdot(\mathrm{kcal/mol})^{1/12}$), is included to prevent unphysical attraction at very short interatomic distances.
-The dispersion term is damped at short ranges to avoid singularities and the double-counting of correlation effects, analogous to the DFT-D3 treatment [18].
+The dispersion term is damped at short ranges to avoid singularities and the double-counting of correlation effects, analogous to the DFT-D3 treatment [grimme_consistent_2010].
 The dimensionless damping factor $s_{\mathrm{disp}}$ is set to 120.
 As demonstrated in Section 1.2, this damped dispersion function provides an excellent fit to the target ALMO-EDA data.
-The electrostatic interaction are modeled following the AMOEBA polarizable framework [11].
+The electrostatic interaction are modeled following the AMOEBA polarizable framework [demerdashAssessingManybodyContributions2017].
 We use atom-centered point charges for the permanent electrostatics and an induced dipole model for polarization:
@@
 For efficient training, $\boldsymbol{\mu}^{\mathrm{ind}}$ is computed via direct matrix inversion, as detailed in Supplementary Section 6.
-The atomic charge $q$ and polarizability alphaalpha used in solving $\mu^{\mathrm{ind}}$ are predicted by the GNN model.
-Finally, the charge-transfer term is modeled using a functional form inspired by the work of Deng et al. [12]:
+The atomic charge $q$ and polarizability $\alpha$ used in solving $\mu^{\mathrm{ind}}$ are predicted by the GNN model.
+Finally, the charge-transfer term is modeled using a functional form inspired by the work of Deng et al. [deng2017estimating]:
 $U_{\mathrm{ct}}^{\mathrm{FF}}(\bm{r};\epsilon^{\mathrm{ct}},\lambda^{\mathrm{ct}},r^{*})=\sum_{i<j}\frac{\epsilon^{\mathrm{ct}}_{ij}}{r_{ij}^{4}}\exp\left(-\left(\frac{\lambda^{\mathrm{ct}}_{ij}r^{*}_{ij}}{r_{ij}}\right)^{3}\right),$ (13)
@@
 Similar to our previous work on ByteFF, atomic charges $q$ are predicted using a Bond Charge Correction (BCC) model to preserve charge conservation.
-Furthermore, a subset of parameters are obtained using a physically-inspired Tkatchenko-Scheffler (TS) scaling model [48]:
+Furthermore, a subset of parameters are obtained using a physically-inspired Tkatchenko-Scheffler (TS) scaling model [tkatchenko_accurate_2009]:
 $C_{6,i}=C_{6,i}^{\mathrm{free}}\left(\frac{V_{i,\mathrm{eff}}}{V_{i,\mathrm{free}}}\right)^{2},\ \alpha_{i}=\alpha_{i}^{\mathrm{free}}\left(\frac{V_{i,\mathrm{eff}}}{V_{i,\mathrm{free}}}\right),\ r_{i}^{*}=r_{i}^{*,\mathrm{free}}\left(\frac{V_{i,\mathrm{eff}}}{V_{i,\mathrm{free}}}\right)^{4/21}.$ (14)
@@
 Once all atom-wise parameters are determined, the pairwise interaction parameters are derived using combination rules.
-We employ the standard Lorentz-Berthelot (LB) rules [36], which use an arithmetic mean for $r_{i}^{*}$ and a geometric mean for all other parameters:
+We employ the standard Lorentz-Berthelot (LB) rules [lorentz_ueber_1881], which use an arithmetic mean for $r_{i}^{*}$ and a geometric mean for all other parameters:
 $r_{ij}^{*}=\frac{r_{i}^{*}+r_{j}^{*}}{2},\quad\epsilon_{ij}=\sqrt{\epsilon_{i}\epsilon_{j}},\quad\lambda_{ij}=\sqrt{\lambda_{i}\lambda_{j}},\quad C_{6,ij}=\sqrt{C_{6,i}C_{6,j}}.$ (15)
@@
 ByteFF-Pol’s training relies on three core datasets: the existing optimization and torsion datasets from our previous work and a newly created dimer dataset.
-Details on the original optimization and torsion datasets are in our previous publication [55].
-For this work, we’ve augmented the optimization dataset by calculating the minimum basis iterative stockholder (MBIS) charges and volumes for each molecule at the PBE0/def2-TZVPD level, which are used to pre-train the model.
+Details on the original optimization and torsion datasets are in our previous publication [zheng2025data].
+For this work, we’ve augmented the optimization dataset by calculating the minimal basis iterative Stockholder (MBIS) charges and volumes for each molecule at the PBE0/def2-TZVPD level, which are used to pre-train the model.
 […]
 To benchmark the performance of ByteFF-Pol, we used five datasets of experimental physical properties.
-The first contains 96 pure molecular liquids with known densities and evaporation enthalpies from the QUBEKit benchmark [23].
-The second and third datasets consist of experimental densities for 595 binary liquids from the NIST ThermoML archive [8] and 2132 pure molecular liquids from the CRC Handbook [35], respectively.
-For electrolyte-relevant systems, we compiled two additional datasets collected from literatures [54]: one with 88 experimental densities and 91 experimental viscosities for pure solvents and lithium electrolytes, and another containing 4896 conductivity measurements for various lithium electrolytes across different compositions and temperatures.
+The first contains 96 pure molecular liquids with known densities and evaporation enthalpies from the QUBEKit benchmark [hortonQUBEKitAutomatingDerivation2019].
+The second and third datasets consist of experimental densities for 595 binary liquids from the NIST ThermoML archive [chirico2003thermoml] and 2132 pure molecular liquids from the CRC Handbook [Lide2004CRC], respectively.
+For electrolyte-relevant systems, we compiled two additional datasets collected from literature [yang2025unified]: one with 88 experimental densities and 91 experimental viscosities for pure solvents and lithium electrolytes, and another containing 4896 conductivity measurements for various lithium electrolytes across different compositions and temperatures.
 The number of results for each force field may differ from that of the benchmark set, owing to parametrization errors or MD simulation failures.
@@
 Our training procedure consists of three stages: pre-training, main training, and fine-tuning. Detailed training techniques are provided in Supplementary Section 7; here, we present a brief overview.
-In the initial pre-training stage, the GNN model is trained to reproduce the bonded parameters from ByteFF [55], along with atomic charges and volumes from MBIS calculations.
+In the initial pre-training stage, the GNN model is trained to reproduce the bonded parameters from ByteFF [zheng2025data], along with atomic charges and volumes from MBIS calculations.
 This stage allows the model to develop atomic hidden states that recognize the chemical environment and provides a physically-grounded starting point for the subsequent training.
@@
 The training targets are the total interaction energies and their decomposed components from the ALMO-EDA calculations.
-Our force field’s energy decomposition scheme is carefully designed to align with the ALMO-EDA scheme, ensuring that the model learns physically meaningful interactions, as detailed in Supplementary Section 7.
+We employ a scaled mean squared error (SMSE) loss for train each energy term:
+$\mathcal{L}_{\mathrm{SMSE}}^{\mathrm{term}}=\frac{1}{N}\sum_{i}s_{i}^{\mathrm{Boltz}}s_{i}^{\mathrm{Force}}\left(U^{\mathrm{FF,inter}}_{\mathrm{term},i}-U^{\mathrm{EDA}}_{\mathrm{term},i}\right)^{2},$ (16)
+where “term” refers to a specific interaction term, i.e., “dispersion”, “electrostatic with Pauli repulsion”, “polarization” and “charge transfer”.
+The two scaling factors, $s_{i}^{\mathrm{Boltz}}$ and $s_{i}^{\mathrm{Force}}$, reduce the weighting of data points with large interaction energies and large interaction forces, respectively.
+This ensures a stable training process and improves the accuracy for conformations that have higher Boltzmann sampling probabilities.
+Detailed definitions of the scaling factors and the decomposed energy terms are provided in Supplementary Section 7.2.
 Finally, in the fine-tuning stage, we update the torsion parameters to ensure they are consistent with the newly trained non-bonded interactions.
-The training target for this stage is the set of torsion scan energy profiles from ByteFF torsion dataset [55].
+The training target for this stage is the set of torsion scan energy profiles from ByteFF torsion dataset [zheng2025data].
 4 Data availability
-The source data for Figure 3 and Figure 4 are provided in the supplementary information. A subset of the training dataset used in this study will be released with the source code.
+A subset of the training dataset and a subset of the validation dataset that is used to reproduce the results in Figure 2 (a) are released with the source code.
+Source Data for Figure 2, 3 and 4 are provided with this paper.
 5 Code availability
-The code for ByteFF-Pol including the training scheme and trained model will be released upon acceptance of this paper.
+The code for ByteFF-Pol including the training scheme, trained model and examples are available at https://github.com/ByteDance-Seed/byteff2.